OpenStudy (anonymous):

how do i solve log (2) x - log (2) 4 = 2? log(2)x/4=2?

OpenStudy (loser66):

log a - log b = log (a/b)

OpenStudy (anonymous):

what do i do with the two since it does not have a log?

OpenStudy (loser66):

take steps, friend, tell me , with my formula, how do you apply to the LHS. what do you get?

OpenStudy (loser66):

oh, I see

OpenStudy (anonymous):

is log(2)x/4=2 the correct set up?

OpenStudy (loser66):

now, 2 = log (2) 1 , right? because 2^1 =2. so you can replace 2 = log(2)1

OpenStudy (loser66):